17 The lower end of a 95% interval estimate of the mean annual salary of persons with an MBA degree five years after graduation is reported to be $81,200.   The interval is based on a sample of 310 persons. The sample standard deviation is $23,356. What is the point estimate of the salary? (Note: The z-score was used in building the interval)
a $85,100
b $84,900
c $84,100
d $83,800
18 In the previous question, what is the upper end of the interval?
a $86,400
b $86,700
c $87,500
d $87,700

I'm doing something wrong because every answer I get is slightly off.
Answer #1

Lower Bound = 81200
std.dev = 23356 , n = 310

LOwer Bound = mean -z *(s/sqrt(n))

81200 = mean - 1.96 *(23356/sqrt(310))
mean = 2600 + 81200

mean = 83800

18)

Upper Bound = 83800 + 1.96 *(23356/sqrt(310))
Upper Bound = 86400